# Likes, Budgets, and Equilibria: Designing Contests for Socially Optimal Advertising

### v1

@@ A.1 Limiting Value of Awareness
$x_{si}(t)$ $=\frac{1}{t}\left[\alpha\sum_{j\in\mathscr{N}_{i}}e_{ji}x_{sj}(t-1)+(1-\alpha)h(b_{si}(k),b_{-si}(k))\right]$
$\quad+\left(1-\frac{1}{t}\right)x_{si}(t-1)$
Let us define the vector $\mathbf{y}_{s}(t)$ of size $n+m$, where the first $n$ entries correspond to the vector $\mathbf{x}_{s}(t)$ and for $\mathbf{y}_{s}(t)]_{n+i}=0$ if $s\neq i$ and $\mathbf{y}_{s}(t)]_{n+s}=1$ if $s=i$. Now consider the matrix $W$, whose element $W_{ij}$ represents the rate of transitioning from state $i$ to $j$:
$W=\begin{bmatrix}-1&\alpha e_{12}&\dots&\alpha e_{1n}&(1-\alpha)h_{11}&\dots&(1-\alpha)h_{m1}\\ \alpha e_{21}&-1&\dots&\alpha e_{2n}&(1-\alpha)h_{12}&\dots&(1-\alpha)h_{m2}\\ \alpha e_{31}&\alpha e_{32}&\dots&\alpha e_{2n}&(1-\alpha)h_{13}&\dots&(1-\alpha)h_{m3}\\ \vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ \alpha e_{n1}&\alpha e_{n2}&\dots&-1&(1-\alpha)h_{1n}&\dots&(1-\alpha)h_{mn}\\ 0&0&\dots&0&0&\dots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&0&0&\dots&0\end{bmatrix}$
@@
$V=\begin{bmatrix}0&\frac{\alpha e_{12}}{W_{1}}&\dots&\frac{\alpha e_{1n}}{W_{1}}&\frac{(1-\alpha)h_{11}}{W_{1}}&\dots&\frac{(1-\alpha)h_{m1}}{W_{1}}\\ \frac{\alpha e_{21}}{W_{2}}&0&\dots&\frac{\alpha e_{2n}}{W_{2}}&\frac{(1-\alpha)h_{12}}{W_{2}}&\dots&\frac{(1-\alpha)h_{m2}}{W_{2}}\\ \vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ \frac{\alpha e_{n1}}{W_{n}}&\frac{\alpha e_{n2}}{W_{n}}&\dots&0&\frac{(1-\alpha)h_{1n}}{W_{n}}&\dots&\frac{(1-\alpha)h_{mn}}{W_{n}}\\ 0&0&\dots&0&1&\dots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&0&0&\dots&1\end{bmatrix}$
@@
Let us suppose that an asymmetric equilibrium exists. Then there exist 2 distinct firms $p\in\mathcal{M}$ and $q\in\mathcal{M}$ such that $\mathbf{b}_{p}\neq\mathbf{b}_{q}$. Because each firm has the same budget, there exist nodes $i$ and $j$ such that $b_{pi}>b_{qi}$ and $b_{qj}>b_{pj}$. For each firm, the KKT conditions require equality of marginal utilities across all active agents:
$c_{i}\frac{\partial h(b_{pi},\mathbf{b}_{-pi})}{\partial x}=c_{j}\frac{\partial h(b_{pj},\mathbf{b}_{-pj})}{\partial x}$
@@ C.2 Network Generation
The network generation process illustrated in Figure˜5 outlines the social network construction described in Section˜5.1. In the following, we provide a detailed explanation of this procedure.

### v2

@@ A.1 Limiting Value of Awareness
$x_{s,i}(t)$ $=\frac{1}{t}\left[\alpha\sum_{j\in\mathscr{N}_{i}}e_{ji}x_{s,j}(t-1)+(1-\alpha)h(b_{s,i}(k),b_{-s,i}(k))\right]$
$\quad+\left(1-\frac{1}{t}\right)x_{s,i}(t-1)$
Let us define the vector $\mathbf{y}_{s}(t)$ of size $n+m$, where the first $n$ entries correspond to the vector $\mathbf{x}_{s}(t)$ and for $\mathbf{y}_{s}(t)]_{n+i}=0$ if $s\neq i$ and $\mathbf{y}_{s}(t)]_{n+s}=1$ if $s=i$. Now consider the $(n+m)\times(n+m)$ interaction matrix $W$, whose elements $W_{ij}$ follow directly from the awareness update in (1): the top-left $n\times n$ block captures awareness propagation between nodes (the first term in (1)), while the last $m$ columns represent direct communication from firms to nodes (the second term in (1)), with all remaining entries set to 0.
$W=\begin{bmatrix}-1&\alpha e_{12}&\dots&\alpha e_{1n}&(1-\alpha)h_{1,1}&\dots&(1-\alpha)h_{m,1}\\ \alpha e_{21}&-1&\dots&\alpha e_{2n}&(1-\alpha)h_{1,2}&\dots&(1-\alpha)h_{m,2}\\ \alpha e_{31}&\alpha e_{32}&\dots&\alpha e_{2n}&(1-\alpha)h_{1,3}&\dots&(1-\alpha)h_{m,3}\\ \vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ \alpha e_{n1}&\alpha e_{n2}&\dots&-1&(1-\alpha)h_{1,n}&\dots&(1-\alpha)h_{m,n}\\ 0&0&\dots&0&0&\dots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&0&0&\dots&0\end{bmatrix}$
@@
$V=\begin{bmatrix}0&\frac{\alpha e_{12}}{W_{1}}&\dots&\frac{\alpha e_{1n}}{W_{1}}&\frac{(1-\alpha)h_{1,1}}{W_{1}}&\dots&\frac{(1-\alpha)h_{m,1}}{W_{1}}\\ \frac{\alpha e_{21}}{W_{2}}&0&\dots&\frac{\alpha e_{2n}}{W_{2}}&\frac{(1-\alpha)h_{1,2}}{W_{2}}&\dots&\frac{(1-\alpha)h_{m,2}}{W_{2}}\\ \vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ \frac{\alpha e_{n1}}{W_{n}}&\frac{\alpha e_{n2}}{W_{n}}&\dots&0&\frac{(1-\alpha)h_{1,n}}{W_{n}}&\dots&\frac{(1-\alpha)h_{m,n}}{W_{n}}\\ 0&0&\dots&0&1&\dots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&0&0&\dots&1\end{bmatrix}$
@@
Let us suppose that an asymmetric equilibrium exists. Then there exist 2 distinct firms $p\in\mathcal{M}$ and $q\in\mathcal{M}$ such that $\mathbf{b}_{p}\neq\mathbf{b}_{q}$. Because each firm has the same budget, there exist nodes $i$ and $j$ such that $b_{p,i}>b_{q,i}$ and $b_{q,j}>b_{p,j}$. For each firm, the KKT conditions require equality of marginal utilities across all active agents:
$c_{i}\frac{\partial h(b_{p,i},\mathbf{b}_{-p,i})}{\partial x}=c_{j}\frac{\partial h(b_{p,j},\mathbf{b}_{-p,j})}{\partial x}$
@@ C.2 Network Generation
The network generation process illustrated in Figure˜4 outlines the social network construction described in Section˜5.1. In the following, we provide a detailed explanation of this procedure.